# JailbreakLens: Interpreting Jailbreak Mechanism in the Lens of Representation and Circuit

### v1

@@ 3 Preliminary
3.1 Large Language Model
An auto-regressive transformer language model $\mathcal{F}$ takes the prompt $x$ as input. The prompt $x$ is comprised with $T$ tokens $\{t_{1},t_{2},...,t_{T}\}$, where each token belongs to a vocabulary set $V$, the model first transforms them into a sequence of token embeddings $\{e_{1},e_{2},...,e_{T}\}$, where each $e_{i}\in\mathbb{R}^{d}$ is transformed by an embedding matrix $\mathbf{W}_{E}\in\mathbb{R}^{|V|\times d}$. These embeddings are deemed as the initial residual stream $h_{i}^{-1}$ for the model. Internally, model $F$ comprises $L$ transformer layers, indexed by $l\in[0,L-1]$. The $l_{th}$ layer gets the information from the residual stream $h_{i}^{l-1}$ as input and write the output of its attention and MLP to this residual stream, updating it to $h_{i}^{l}$. This process can be presented as $h_{i}^{l}=h_{i}^{l-1}+m_{i}^{l}+a_{i}^{l}$, where $m_{i}^{l}$ and $a_{i}^{l}$ are the outputs from the attention component and MLP component, respectively. For simplicity, we omit the layer normalization of each module.
After finishing the last layer, we obtain the logit values of the last token over the vocabulary space $v_{T}\in\mathbb{R}^{|V|}$ using an unembedding matrix $\mathbf{W}_{U}\in\mathbb{R}^{d\times|V|}$ where $v_{T}=\mathbf{W}_{U}(h_{T}^{L-1})$ from which we can sample a new token. This process continues iteratively, generating one token at a time until the model encounters the end-of-sequence token, denoted as $w_{eos}$, at which point the generation process halts.
3.2 Interpretability Techniques
@@
Probing. The probing technique [34, 35, 36] aims to reveal what information the model has learned and stored in the intermediate representation, i.e., $h_{i}^{l}$, in different model layers. At each layer, the output generated by the model’s attention and MLP modules will update the residual stream $h_{i}^{l}$, and these updates may encode some semantic, syntactic, and other information of the input. The probe classifier $p$ predicts the feature $z$, such as a part-of-speech tag, or semantic and syntactic information including emotion, in the intermediate representation by training a supervised model, i.e., $p:h_{i}^{l}\rightarrow z$, to determine how much input-specific information the model stores in different layers. By probing the representation layer by layer, we can understand how the model encodes and extracts information during the generation process, revealing the role of each layer in representation and information transfer.
@@
Logitlen. Logitlen [37, 38, 39], interprets the information encoded in the hidden states of each layer by mapping them back to the vocabulary space. In particular, logitlen can not only be used to explain the final output representation $h_{i}^{l}$ of each transformer block but can also be refined to interpret the output of each attention head (i.e., $a_{i}^{l}$) or MLP module (i.e., $m_{i}^{l}$). Specifically, logitlen maps the hidden state of each component to the vocabulary distribution through the unembedding matrix $\mathbf{W}_{U}$, so that we can observe the changes in vocabulary distribution layer by layer during the generation process and analyze How the model processes and understands input information at different layers. The logitlen is worked as $p=\mathbf{W}_{U}(h)$, where $h$ denotes the hidden states to be interpreted and $p$ denotes the logit for each token in the vocabulary set $V$.
4 JailbreakLens: Dual-Perspective Interpretation Framework
@@
In this section, we propose a dual-perspective interpretation framework, named JailbreakLens, illustrated in Fig. 1, which investigates jailbreak mechanisms from both representation and circuit perspectives. Representation level analysis examines how different types of prompts, i.e., safe prompts, harmful prompts, and jailbreak prompts, are mapped within the latent space, helping to identify vulnerable areas of the model at the semantic level. For example, we can observe whether jailbreak prompts are more likely to be mapped to specific semantic regions associated with safe prompts, triggering uncontrolled generation behavior. Circuit level analysis deeply studies which model components, including attention heads and MLP layers in each transformer layer, play key roles in model safeguard and how jailbreaks affect their behaviors to cause the model to produce unintended outputs. Such dual-perspective analysis can not only identify how jailbreaks deceive the model’s semantic understanding of harmfulness but also locate which model components contribute highly to these vulnerabilities.
@@
  • RQ3. How do internal representations and key model components evolve throughout the token generation process during the jailbreak attack? While two above questions reveal how jailbreak prompts affect the model’s harmfulness perception and identify the critical components involved, understanding how these disruptions progress through each token generation step is essential. Tracking the dynamic changes in representations and key components across the generation process allows us to observe how jailbreak attacks gradually undermine the model’s safeguards, leading to harmful outputs, in Sec. 4.3.
@@
4.1 Representation Probing
For representation level analysis, our goal is to analyze how the jailbreak prompts change the model’s perception of safe and harmful content in the representation space, thereby achieving the purpose of bypassing security mechanisms. Specifically, representation level interpretation is divided into three steps, i.e., evaluating the model’s harmfulness perception, detecting harmfulness within jailbreak prompts, and analyzing the model’s response tone across layers.
To evaluate the model’s harmfulness perception, we train a probing classifier, denoted as $\mathcal{P}$, to assess whether a model can correctly distinguish the harmful prompts from the safe ones. The safety probe $\mathcal{P}$ is trained on a binary classification task with the dataset $\mathcal{D}=(X_{+},X_{-})$, consisting of safe prompts (i.e., $X_{+}$ labeled as +1 ) and harmful prompts (i.e., $X_{-}$ labeled as -1). We firstly obtain model representations from each layer on dataset $\mathcal{D}$, and utilize this representation, i.e., $A_{+}^{l}$ as the representation on safe prompts on $l_{th}$ layer and $A_{-}^{l}$ as the that on harmful prompts, as the input of probe $\mathcal{P}$. The training dataset for probes is denoted as $\mathcal{D_{A}}=(A_{+},A_{-})$, with each sample $a_{i}\in\mathbb{R}_{d}$, where $a_{i}$ is an individual sample in $D_{A}$ and $d$, is the dimension of the model’s hidden states. Specifically, we train the safety probes with different architectures, including neural network based, clustering based, and dimension reduction based.
Moreover, we can obtain a safety direction vector, denoted as $v_{d}$, via the well-trained probe. For instance, safety direction vector of cluster probe can be calculated the difference between the center of harmful representation cluster and the center of safe representation cluster. To evaluate the semantic meaning of the direction vector, we project it into the vocabulary space via logitlen to verify whether they align with the expected ‘safe-harmful’ direction.
To detect harmfulness within jailbreak prompts, we classify the jailbreak representations with well-trained safety probes. If the well-trained probe classifies a representation generated by a jailbreak prompt as safe, suggesting the attack has successfully misled the model. Conversely, if the probe identifies the representation as harmful, the model is still able to recognize the malicious intent of the prompt, indicating the deception was unsuccessful.
@@ 4.1 Representation Probing
4.2 Model Behavior Circuit
@@
While representation analysis helps uncover how the model’s perception of harmfulness shifts during jailbreak attacks, we further delve into the internal circuits to analyze the underlying cause of the representation level changes, specifically focusing on attention heads and MLP layers in the transformer blocks. For circuit level analysis, our goal is to identify model components that contribute significantly to harmful perception and explore how jailbreak cues influence the behavior of these key components.
@@
To identify the model components that contribute highly to generation safety, We employ logit attribution techniques to quantify each component’s specific contribution to the output. By direct logit attribution $\mathcal{F}c(x)W{U[:,w]}$, we can measure the impact of a specific model component $\mathcal{F}_{c}$ on the predicted token $w$. Next, in order to further quantify the model’s preference for safe or harmful output, we need to select two target tokens, namely a token $w_{+}$ representing a safe response and a token $w_{-}$ representing a harmful response. To this end, we first obtain the safety direction vector $v$ from the representation layer analysis, project it into the vocabulary space with logitlen to select the most representative positive attribution target token, that is:
$w_{+}=\arg\max_{w\in V}\mathbf{W}_{U}(v),$ (1)
where $V$ denotes the vocabulary set, $\mathbf{W}_{U}$ denotes the unembedding matrix of the model. And we use the reverse vector of the safety direction vector, i.e., $-v$, to obtain the negative attribution target token, i.e., $w_{-}$.
@@
After determining the target tokens $w_{+}$ and $w_{-}$ for attribution analysis, we introduce a ”refusal score” (named $rs$) to evaluate each model component’s role in generating a safe response, which is defined as the logit difference between the positive target token and negative target token. Mathematically, the refusal score is defined in Eq. (2):
$rs=F^{c}(\mathbf{x})W_{U[:,w_{-}]}-F_{c}(\mathbf{x})W_{U[:,w_{+}]},$ (2)
@@
where $F^{c}$ denotes the model component under measurement, $w_{-}$ denotes positive target token, and $w_{+}$ denotes negative target token.
@@
The workflow of computing the contribution of circuits on generation safety is summarized in Alg. 1. Firstly, we obtain the two attribution target tokens, i.e., $w_{+}$ and $w_{-}$, for model components in each transformer layer. Then we qualify the contribution of each component $\mathcal{F}_{c}$, such as an individual attention head or MLP layer, on prompt set $X_{t}$ via calculating the refusal score with the tokens we identified. Specifically, we use the harmful dataset $X_{-}$ as the test set $X_{t}$ in Alg. 1 to identify the key components that contribute to refusal (named refusal signal component, $S_{-}$) and with the safe dataset $X_{+}$ as the test set $X_{t}$ to identify the key components that contribute to affirmation (named refusal signal component, $S_{+}$). The components with highest $rs$ would be identified as key components on safety. Then we investigate how jailbreak prompts influence these components by comparing the activation of these components on safe prompts, harmful prompts, and jailbreak prompts.
Algorithm 1 Component Importance Calculation
@@
Input: Safety direction vector set $\{v_{1},v_{2},...,v_{L}\}$, testing prompt set $X_{t}$, component $\mathcal{F}_{c}$ in $l$ layer.
Output: Causal effects for $\mathcal{F}_{c}$: $rs_{c}$.
$w_{+}=\arg\max_{w\in V}\mathbf{W}_{U}(v_{l})$triangleright positive target token
$w_{-}=\arg\max_{w\in V}\mathbf{W}_{U}(-v_{l})$triangleright negative target token
for $(X_{t}^{(i)}$) in $X_{t}$ do
     $prob_{w_{-}}\leftarrow\mathcal{F}_{c}(X_{t}^{(i)})W_{U[:,w_{-}]}$
     $prob_{w_{+}}\leftarrow\mathcal{F}_{c}(X_{t}^{(i)})W_{U[:,w_{+}]}$
     ${rs_{c}^{(i)}}\leftarrow prob_{w_{-}}-prob_{w_{+}}$triangleright refusal score
end for
Return: ${rs_{c}}=\frac{\sum_{i=1}^{|X_{t}|}rs_{c}^{(i)}}{|\mathbf{X_{t}}|}$triangleright averaged effect
4.3 Interpretation Aggregation
@@
In order to explain the jailbreak mechanism more accurately, we extend our analysis to the dynamic evolution of the model representation and the activation of key components throughout the token generation sequence in response to jailbreak prompts.
For representation analysis, we evaluate the harmfulness level of the representation from each token generation. To measure this, we use the well-trained safety probe $\mathcal{P}$ to predict representations from each generated token. Specifically, if the probing logit is positive, the larger the value, the higher the affirmation level of the representation. On the contrary, if the probing logit is negative, the smaller the value, the higher the refusal of the representation. For the circuit dynamic analysis, we track the evolution of signal components, i.e., $S_{+}$ and $S_{-}$, via observing their activation during the generation of each token, specifically whether each component is enhanced or suppressed. The workflow of tracking the evolutions is summarized in Alg. 2.
Algorithm 2 Tracking the Evolution of Representation and Circuit
@@
Input: Jailbreak prompt $x$, model $\mathcal{F}$ with layer $l$’s representation under measurement, signal components ($S_{+},S_{-}$), well-trained probe $\mathcal{P}$, and signal tokens ($w_{+},w_{-}$).
Output: Evolution for representations and key components: ($E_{R},E_{C}$).
$Y\leftarrow[]$triangleright start with an empty generated sequence
$i\leftarrow 0$triangleright start generating from the first token
while $y_{i}\neq w_{eos}$ do $\triangleright$ continue until end-of-sequence token $w_{eos}$ is generated
     $A_{l}^{(i)}\leftarrow\mathcal{F}_{l}(x+Y)$
     $E_{R}^{(i)}\leftarrow\mathcal{P}(A_{l}^{(i)})$

### v2

@@ 3. Preliminary
3.1. Large Language Model
An auto-regressive language model $\mathcal{F}$ takes the prompt $x$ as input, which consisists of $T$ tokens $\{t_{1},t_{2},...,t_{T}\}$ and each token belongs to a vocabulary set $V$. The model first transforms $x$ into embeddings $\{e_{1},e_{2},...,e_{T}\}$, where each $e_{i}\in\mathbb{R}^{d}$ is produced by an embedding matrix $\mathbf{W}_{E}\in\mathbb{R}^{|V|\times d}$. These embeddings are the initial model residual stream $h_{i}^{-1}$. Model $F$ comprises $L$ transformer layers, indexed by $l\in[0,L-1]$. The $l_{th}$ layer processes the residual stream $h_{i}^{l-1}$ and updates it to $h_{i}^{l}$ through attention and MLP. This process can be presented as $h_{i}^{l}=h_{i}^{l-1}+m_{i}^{l}+a_{i}^{l}$, where $m_{i}^{l}$ and $a_{i}^{l}$ are the outputs from the attention component and MLP component, respectively. For simplicity, we omit the layer normalization of each module.
After the final layer, the model computes logit values for the last token across the vocabulary space $v_{T}\in\mathbb{R}^{|V|}$ using an unembedding matrix $\mathbf{W}_{U}\in\mathbb{R}^{d\times|V|}$ where $v_{T}=(h_{T}^{L-1})^{\top}\mathbf{W}_{U}$. From these logits, the next token is sampled. Generation continues iteratively until reaching the end-of-sequence token $w_{eos}$.
3.2. Interpretability Techniques
@@
Probing. The probing technique (Belinkov, 2022; Gurnee and Tegmark, 2023; Marks and Tegmark, 2023) aims to reveal what information the model has learned and stored in the intermediate representation, i.e., $h_{i}^{l}$, in different model layers. At each layer, the output generated by the model’s attention and MLP modules will update the residual stream $h_{i}^{l}$, and these updates may encode some semantic, syntactic, and other information of the input. The probe classifier $p$ predicts the feature $z$, such as a part-of-speech tag, or semantic and syntactic information including emotion, in the intermediate representation by training a supervised model, i.e., $p:h_{i}^{l}\rightarrow z$, to determine how much input-specific information the model stores in different layers. By probing the representation layer by layer, we can understand how the model encodes and extracts information during the generation process, revealing the role of each layer in representation and information transfer.
@@
Logitlen. Logitlen (les, [n. d.]; Sakarvadia et al., 2023; Belrose et al., 2023) interprets the information encoded in the hidden states of each layer by mapping them back to the vocabulary space. In particular, logitlen can not only be used to explain the final output representation $h_{i}^{l}$ of each transformer block but can also be refined to interpret the output of each attention head (i.e., $a_{i}^{l}$) or MLP module (i.e., $m_{i}^{l}$). Specifically, logitlen maps the hidden state of each component to the vocabulary distribution through the unembedding matrix $\mathbf{W}_{U}$, so that we can observe the changes in vocabulary distribution layer by layer during the generation process and analyze how the model processes and understands input information at different layers. The logitlen is worked as $p=h^{\top}\mathbf{W}_{U}$, where $h$ denotes the hidden states to be interpreted and $p$ denotes the logit for each token in the vocabulary set $V$.
4. JailbreakLens: Dual-Perspective Interpretation Framework
@@
In this section, we propose a dual-perspective interpretation framework, named JailbreakLens, illustrated in Fig. 1, which investigates jailbreak mechanisms from both representation and circuit perspectives. Representation level analysis examines how different types of prompts, i.e., safe prompts, harmful prompts, and jailbreak prompts, are mapped within the latent space, helping to identify vulnerable areas of the model at the semantic level. For example, we can observe whether jailbreak prompts are more likely to be mapped to specific semantic regions associated with safe prompts, triggering uncontrolled generation behavior. Circuit level analysis deeply studies which model components, including attention heads and MLP layers in each transformer layer, which play key roles in model safeguard and how jailbreaks affect their behaviors to cause the model to produce unintended outputs. Such dual-perspective analysis can not only identify how jailbreaks deceive the model’s semantic understanding of harmfulness but also locate which model components contribute highly to these vulnerabilities.
@@
  • RQ3. How do internal representations and key model components evolve throughout the token generation process during the jailbreak attack? While above questions reveal how jailbreaks affect model’s harmfulness perception and identify the critical circuits involved, a complete understanding requires analyzing how these effects evolve throughout the entire generation process. We conduct two complementary analyses, i.e., correlation analysis between representational deception and key circuit activation shifts which quantifies the alignment between semantic manipulation and functional disruption) and dynamic analysis that traces the token-by-token evolution of both representation safety perception and key circuit activation.
@@
4.1. Representation Probing
For representation level analysis, our goal is to analyze how jailbreak prompts change the model’s perception of safe and harmful content in representation space, thereby achieving the purpose of bypassing security mechanisms. Specifically, representation level interpretation is divided into three steps, i.e., evaluating model’s harmfulness perception ability, probing deceptive harmfulness of jailbreak prompts, and analyzing model’s response tone across each model layer.
To evaluate model’s harmfulness perception ability, we train a probing classifier, denoted as $\mathcal{P}$, to assess whether a model can correctly distinguish harmful prompts from the safe ones. The safety probe $\mathcal{P}$ is trained on a binary classification task with dataset $\mathcal{D}=(X_{+},X_{-})$, consisting of safe prompts (i.e., $X_{+}$ labeled as +1 ) and harmful prompts (i.e., $X_{-}$ labeled as -1). We extract model representations from each layer on dataset $\mathcal{D}$, and utilize this representation, i.e., $A_{+}^{l}$ as the representation on safe prompts on $l_{th}$ layer and $A_{-}^{l}$ as the that on harmful prompts, as the input of probe $\mathcal{P}$. The training dataset for probes is denoted as $\mathcal{D_{A}}=(A_{+},A_{-})$, with each sample $a_{i}\in\mathbb{R}_{d}$, where $a_{i}$ is an individual sample in $D_{A}$, and $d$ is the dimension of the hidden states. Specifically, we train the safety probes with different architectures, including neural network based, clustering based, and dimension reduction based.
Moreover, we can obtain a safety direction vector, denoted as $v_{d}$, via the well-trained probe $\mathcal{P}$. For instance, safety direction vector of cluster probe can be calculated the difference between the center of harmful representation cluster and the center of safe representation cluster. To evaluate the semantic meaning of direction vector, we project it into the vocabulary space via logitlen to verify whether they align with the expected ‘safe-harmful’ direction.
To detect harmfulness within jailbreak prompts, we classify the jailbreak representations with well-trained safety probes. If the well-trained probe classifies a representation generated by a jailbreak prompt as safe, suggesting the attack has successfully misled the model. Conversely, if the probe identifies the representation as harmful, the model is still able to recognize the malicious intent of the prompt, indicating the deception is unsuccessful.
@@ 4.1. Representation Probing
4.2. Model Behavior Circuit
@@
While representation analysis helps uncover how the model’s perception of harmfulness shifts during jailbreak attacks, we further delve into internal model circuits to analyze the underlying cause of representation level changes, specifically focusing on attention heads and MLP layers in transformer blocks. For circuit level analysis, our goal is to identify model components that contribute significantly to harmfulness perception and explore how jailbreak prompts influence the behavior of these key components.
@@
To identify the model components that contribute highly to generation safety, we employ logit attribution techniques to quantify each component’s specific contribution to the output. By direct logit attribution $\mathcal{F}_{c}(x)W_{U[:,w]}$, we can measure the impact of a specific model component $\mathcal{F}_{c}$ on the predicted token $w$. Next, in order to further quantify the model’s preference for safe or harmful output, we need to select two target tokens, namely a token $w_{+}$ representing a safe response and a token $w_{-}$ representing a harmful response. To this end, we first obtain the safety direction vector $v$ from the representation layer analysis, project it into the vocabulary space with logitlen to select the most representative attribution target token, that is:
(1) $w_{+}$ $=\arg\max_{w\in V}\langle v,\mathbf{W}_{U[:,w]}\rangle,$
(2) $w_{-}$ $=\arg\max_{w\in V}\langle-v,\mathbf{W}_{U[:,w]}\rangle.$
[…]
@@
After determining the target tokens $w_{+}$ and $w_{-}$ for attribution analysis, we introduce a ”refusal score” ($rs$) as the logit difference between positive and negative target tokens to assess each model component’s role in generating safe responses. Mathematically, $rs$ is defined in Eq. (3):
(3) $rs=\mathcal{F}_{c}(\mathbf{x})W_{U[:,w_{-}]}-\mathcal{F}_{c}(\mathbf{x})W_{U[:,w_{+}]},$
@@
where $\mathcal{F}_{c}$ denotes the model component under measurement, $w_{+}$ denotes positive target token, and $w_{-}$ denotes negative target token.
@@
The workflow of computing the contribution of circuits on generation safety is summarized in Alg. 1. Firstly, we obtain the two attribution target tokens, i.e., $w_{+}$ and $w_{-}$, for model components in each transformer layer. Then we qualify the contribution of each component $\mathcal{F}_{c}$, such as an individual attention head or MLP layer, on prompt set $X_{t}$ via calculating the refusal score with the tokens we identified. Specifically, we use the harmful dataset $X_{-}$ as the test set $X_{t}$ in Alg. 1 to identify the key components that contribute to refusal (named refusal signal component, $S_{-}$) and with the safe dataset $X_{+}$ as the test set $X_{t}$ to identify the key components that contribute to affirmation (named affirmation signal component, $S_{+}$). The components with highest $rs$ would be identified as key circuits on safety. Then we investigate how jailbreak prompts influence these components by comparing the activation of these components on safe prompts, harmful prompts, and jailbreak prompts.
@@
Input: Safety direction vector set $\{v_{1},v_{2},...,v_{L}\}$, testing prompt set $X_{t}$, component $\mathcal{F}_{c}$ in $l$ layer
Output: Causal effects for $\mathcal{F}_{c}$: $rs_{c}$
1 $w_{+}=\arg\max_{w\in V}\langle v_{l},\mathbf{W}_{U[:,w]}\rangle$
2 $w_{-}=\arg\max_{w\in V}\langle-v_{l},\mathbf{W}_{U[:,w]}\rangle$
3 for $(X_{t}^{(i)})$ in $X_{t}$ do
4       $prob_{w_{-}}\leftarrow\mathcal{F}_{c}(X_{t}^{(i)})\mathbf{W}_{U[:,w_{-}]}$
5       $prob_{w_{+}}\leftarrow\mathcal{F}_{c}(X_{t}^{(i)})\mathbf{W}_{U[:,w_{+}]}$
       ${rs_{c}^{(i)}}\leftarrow prob_{w_{-}}-prob_{w_{+}}$ // refusal score
6
7 end for
return ${rs_{c}}=\frac{\sum_{i=1}^{|X_{t}|}rs_{c}^{(i)}}{|\mathbf{X_{t}}|}$ // averaged effect
Algorithm 1 Circuit Importance Calculation
4.3. Interpretation Aggregation
@@
To further bridge the gap between representation-level and circuit-level interpretations, we analyze the quantitive correlation between representation performance and key circuits activations of jailbreaks. Moreover, we extend our analysis to dynamic evolution of model representation and activation of key circuits throughout the token generation sequence in response to jailbreak prompts.
For quantitive correlation analysis, we compute the Pearson correlation coefficient between (1) the deception degree of representation harmfulness under jailbreaks, and (2) activation shifts of safety-related circuit. Their precise definitions are detailed below.
Representation deception measures the probability that a jailbreak prompt is classified as harmful by a pre-trained safety probe $\mathcal{P}$, i.e., defined as $\mathcal{P}(\text{safe}|x)\in[0,1]$. It reflects the model’s internal perception of the prompt’s harmfulness, with higher values indicating stronger deception of malicious intent.
Activation shift quantifies how the activation of key safety-related components (i.e., $S_{+}$ and $S_{-}$) changes in response to $x$ compared to the average behavior on benign ($X_{+}$) and harmful ($X_{-}$) prompts. Formally, we define the activation shift as:
(4) $\Delta A=[S_{+}(x)-S_{+}(X_{+})]+[S_{-}(X_{-})-S_{-}(x)],$
where $x$ is the jailbreak prompt under analysis, $X_{+}$ is the harmless prompt set, $X_{-}$ is the harmful prompt set, $S(~{}\cdot~{})$ is the prompts’ activation on the $S$ components. Constant $S_{+}(X_{+})$ and $S_{-}(x_{-})$ represents the average activations of safe/harmful prompt sets on respective components, providing a reference model state under normal conditions (i.e., when processing safe and harmful prompts). Since $S_{+}(X_{+})$ and $S_{-}(X_{-})$ are constant and our analysis focuses on linear correlation, we can think of it as the relationship between representation toxicity and the activation shift $\Delta A=S_{+}(x)-S_{-}(x)$ which describes the deviation in behavior caused by the jailbreak prompt. The constant term serves as a baseline, used to measure the impact of a specific prompt on the model’s behavior.
For dynamic analysis, we track the evolution of representation deception level and key circuit activation shift degree throughout the whole generation sequence.
To track the representation deception degree, we evaluate the harmfulness level of each generated token’s representation using the well-trained safety probe $\mathcal{P}$. Specifically, a positive probing logit suggests that the probe considers the token harmless, and the larger the value, the higher the affirmation level of the representation. In contrast, a negative probing logit indicates a higher likelihood of the token being harmful, with smaller values corresponding to a stronger refusal.
For the circuit dynamic analysis, we track the evolution of these key signal components (identifed in Sec. 4.2, i.e., $S_{+}$ and $S_{-}$) by observing their activation during the generation of each token. Specifically, we analyze whether each component is enhanced or suppressed. This stage aims to validate and reinforce the consistency between representation-level and the circuit-level performance. By comparing how the safety probe’s evaluation of token harmfulness aligns with the activation patterns of the underlying circuits, we gain deeper insight into how both semantic-level representations and function-level circuit activations contribute to the model’s overall behavior. The workflow for tracking these evolutions is summarized in Alg. 2.
@@
Input: Jailbreak prompt $x$, model $\mathcal{F}$ with layer $l$’s representation under measurement, signal components ($S_{+},S_{-}$), well-trained probe $\mathcal{P}$, and signal tokens ($w_{+},w_{-}$)
Output: Evolution for representations and circuits: ($E_{R},E_{C}$)
$Y\leftarrow[]$ // start with an empty generated sequence
$i\leftarrow 0$ // start generating from the first token
while $y_{i}\neq w_{eos}$ // continue until $w_{eos}$ is generated
1 do
2       $A_{l}^{(i)}\leftarrow\mathcal{F}_{l}(x+Y)$
3       $E_{R}^{(i)}\leftarrow\mathcal{P}(A_{l}^{(i)})$